jigglypuff314 (jigglypuff314):

Use imclicit differentiation to find y' xe^y - 10x + 3y = 0 I'm having trouble with understand how to do this... :/

OpenStudy (kainui):

Take your best guess and I'll help you out. I love this stuff!

jigglypuff314 (jigglypuff314):

-.-" *brain gets fried* erm... is derivative xe^y -> e^y ???

OpenStudy (kainui):

The example I like to show is: \[y=x^2\] algebra doesn't change how one variable changes with respect to the other. \[y^{1/2}=x\] take the derivative of both sides with respect to x\[\frac{1}{2}y^{-1/2}*\frac{dy}{dx}=1x^0*\frac{dx}{dx}\] So all I did there was the chain and power rules. Notice how that simplifies to: \[\frac{dy}{dx}=2y^{1/2}=2x\] which is what we knew all along. If any of that is confusing, ask!

jigglypuff314 (jigglypuff314):

...? how does that apply to my problem? :/

OpenStudy (kainui):

I am showing you how implicit differentiation works.

jigglypuff314 (jigglypuff314):

*nods* but idk how derivative of xe^y works

OpenStudy (kainui):

You just need to use the product rule.

jigglypuff314 (jigglypuff314):

??? xy'e^y + e^y ???

OpenStudy (kainui):

\[\frac{d}{dx}(xe^y)=\frac{d}{dx}(x)e^y+x \frac{d}{dx}(e^y)\]

OpenStudy (kainui):

Yep you got it. =)

jigglypuff314 (jigglypuff314):

y' = (10 - e^y) / (xe^y + 3) ???

OpenStudy (anonymous):

Is that your final answer?

OpenStudy (anonymous):

Wait, yay you're right jiggly ^.^

OpenStudy (anonymous):

Gj!

jigglypuff314 (jigglypuff314):

YAYYYY THANK YOU!!! :D

OpenStudy (anonymous):

This stuff is awesome, as kai mentioned haha.

jigglypuff314 (jigglypuff314):

.-. your smartie panties

OpenStudy (♪chibiterasu):

tee-hee, panties .-.

OpenStudy (anonymous):

I hate the d/dx notation though haha, I'm more of y' kind of guy :)

jigglypuff314 (jigglypuff314):

yrsh easier to type :P

OpenStudy (kainui):

I guess I just think it's more "true" to what's going on. Whatever works lol.